# Large Language Model Enabled Multi-Task Physical Layer Network

### v1

@@ Abstract
The recent advance of Artificial Intelligence (AI) is continuously reshaping the future 6G wireless communications. Recently, the development of Large Language Models (LLMs) offers a promising approach to effectively improve the performance and generalization for different physical layer tasks. However, most existing works finetune dedicated LLM networks for a single wireless communication task separately. Thus performing diverse physical layer tasks introduces extremely high training resources, memory usage, and deployment costs. To solve the problem, we propose a LLM-enabled multi-task physical layer network to unify multiple tasks with a single LLM. Specifically, we first propose a multi-task LLM framework, which finetunes LLM to perform multi-user precoding, signal detection and channel prediction simultaneously. Besides, multi-task instruction module, input encoders, as well as output decoders, are elaborately designed to distinguish multiple tasks and adapted the features of different formats of wireless data for the features of LLM. Numerical simulations are also displayed to verify the effectiveness of the proposed method.
@@ I Introduction
The deep integration of Artificial Intelligence (AI) with wireless communications is one of the key features of Sixth-Generation (6G) communications [1]. With the development of 6G communications, the application of AI becomes important and essential to manage increasing complexity and user demands, and thereby supports more intelligent services and applications. In physical layer communications, owing to strong learning ability from data, AI especially deep learning (DL), has gained significant attention [2] and demonstrated great potential in a wide range of tasks, including CSI feedback [3], channel estimation [4] and prediction [5], signal detection [6], etc. Moreover, significant efforts have been made by 3rd Generation Partnership Project (3GPP) to standardize DL in wireless networks recently [7].
@@
Despite the significant progress, existing DL empowered methods still face some fundamental issues limiting their application in practical scenarios. First, the capability of DL is constrained by the size of networks. With the increasingly high-dimension channel, complex network and rapidly changing environments, existing methods struggle to accurately model the problem and capture features from a global perspective. Secondly, the deep learning-based methods are often sensitive to different wireless environments, and result in poor robustness and generalization, requiring retraining when the channel distribution changes. Recently, Large language models (LLMs) [8, 9, 10] have revolutionized AI domains and attracted considerable interest from both academia and industry, which offer an entirely new paradigm for solving the challenges of existing DL methods. Unlike previous DL algorithms, these models with a huge amount of parameters, possess the ability to capture universal knowledge and demonstrate impressive language understanding and generation capabilities to various tasks in different domains.
Therefore, explosive attention has explored the potential transformative impact of LLMs on wireless communications [11, 12, 13, 14, 15]. Moreover, a few works have focused on opening up new possibilities for improving the performance of different applications in physical layer communications. The authors in [16] propose a LLM-empowered channel prediction method to achieve improved predictive capability and generalization ability. [17] also unleashes the strength of LLMs for time series forecasting to improve the robustness of beam prediction. Besides, in [18] LLM is adopted to perform resource allocation by a few-shot learning-based approach, where the channel gains and the corresponding transmit power strategies are provided as references in the prompt of LLM, confirming the applicability and feasibility of LLM. However, most existing works finetune dedicated LLM networks for a single wireless communication task. Given the parameter size of LLM, designing and training dedicated models for each task separately leads to extremely high computational complexity, memory usage, and deployment costs.
Fortunately, thanks to the excellent language understanding and generation capabilities of LLMs, it is promising to input different task requirements to LLMs with natural language and thus to facilitate performing multiple tasks with one network. Therefore, in this work, we make the initial attempt to propose a LLM-enabled multi-task physical layer network to unify multiple tasks with a single LLM. To be specific, firstly, leveraging the global feature extraction and reasoning ability of LLMs, a multi-task LLM framework is proposed, which finetunes LLM to perform multi-user precoding, signal detection and channel prediction simultaneously. Secondly, to cope with different tasks and different data formats of the tasks, we elaborately design multi-task instructions as prompts of LLM; besides, to bridge the gap between the feature of wireless data and that of the LLM, task-specific encoders and decoders are proposed to adapt the text-based pre-trained LLM. Last but not least, simulation results verified the effectiveness of the proposed method, which achieves comparable performance with dedicated designed network for each task.
@@ II Systems Model
We consider a multi-user (MU) multiple input single output (MISO)- orthogonal frequency division multiplexing (OFDM) system, where a base station (BS) simultaneously serves $K$ single-antenna moving users with $K$ RF chains. The BS is equipped with a uniform planar array (UPA) with $N_{\rm T}=N_{h}\times N_{v}$ antennas, where $N_{h}$ and $N_{v}$ represent the number of antennas in the horizontal and vertical directions, respectively. To design a multi-task LLM for the BS, we select three typical tasks in physical layer communications in different scenarios for instance, namely downlink multi-user precoding, uplink signal detection and channel prediction for moving conditions, respectively. Then, we will briefly illustrate the system models and formulate the tasks sequentially. Note that we use $a$ for downlink transmission while $\bar{a}$ for uplink transmission.
@@ II-A Multi-user Precoding
For downlink transmission scenario, the channel between user $k$ and the BS at the $m$-th subcarrier is denoted as ${\bf h}_{k}^{m}\in\mathbb{C}^{N_{t}\times 1},m=1,2,\cdots,M$. The received signal at user $k$ at the $m$-th subcarrier is given by
$y_{k}^{m}=\mathbf{h}_{k}^{mH}\sum_{k^{\prime}=1}^{K}\mathbf{w}_{k^{\prime}}^{m}x_{k^{\prime}}^{m}+n_{k}^{m},$ (1)
where $\mathbf{w}_{k}^{m}$ represents the beamforming vector for user $k$, $x_{k}^{m}\sim\mathcal{CN}(0,1)$ is the transmitted symbol from the BS to user $k$, and $n_{k}^{m}\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive Gaussian white noise (AWGN) with zero mean and variance $\sigma^{2}$.
Multi-user precoding aims to optimize the transmit beamformers such that the system sum-rate is maximized subject to a total power constraint due to the BS power budget. For simplicity, we design the beamformers based on the channel of the central carrier-frequency ${\bf h}_{k}$ and the problem is mathematically formulated as
@@
where $\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\cdots,\mathbf{w}_{K}]$ is a set of beamforming vectors and $P_{\rm max}$ is the power budget. Besides, $\gamma_{k}$ represents the received signal-to-interference-plus-noise ratio (SINR) at user $k$, which is written as
As pointed out in [19] that the optimal downlink beamforming vectors for problem P1 follows the structure as
where ${\lambda_{k}}$ is a positive parameter and $\sum_{k=1}^{K}{\lambda_{k}}=\sum_{k=1}^{K}p_{k}=P_{\rm max}$. The solution structure in (4) provides the required expert knowledge for the beamforming design in problem P1. Therefore, the network is only required to learn the key features $\mathbf{\lambda}=[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$ and $\mathbf{p}=[p_{1},p_{2},\cdots,p_{K}]$.
@@ II-B Signal Detection
During uplink transmission, the BS receiver antennas simultaneously receive the transmitted symbols from the $K$ users. Denote the transmitted symbol vector from four users at the $m$-th subcarrier as $\bar{\mathbf{x}}^{m}=[\bar{x}^{m}_{1},\bar{x}^{m}_{2},\cdots,\bar{x}^{m}_{K}]\in\mathbb{C}^{K\times 1}$. Each element is drawn from the $P$-QAM constellation, and then transmitted over channel. The received signal $\bar{\mathbf{y}}^{m}\in\mathbb{C}^{N_{\rm T}\times 1}$ is given by
@@
where $\bar{\mathbf{H}}^{m}=[\bar{\mathbf{h}}_{1}^{m},\bar{\mathbf{h}}_{2}^{m}.\cdots,\bar{\mathbf{h}}_{K}^{m}]$ is the uplink channel of the $K$ users and $\bar{\mathbf{n}}^{m}\sim\mathcal{CN}(0,\bar{\sigma}^{2}I_{N_{\rm T}})$ is the AWGN at the $m$-th subcarrier.
BS requires to recover the signals $\bar{\mathbf{x}}^{m}$ from the received signal $\bar{\mathbf{y}}^{m}$ given $\bar{\mathbf{H}}^{m}$. In this work, we adopt the minimum mean squared error (MMSE) estimator and the problem can be formulated as
@@
s.t. $\hat{\bar{\mathbf{x}}}^{m}=f_{\Omega_{\rm DEC}}(\bar{\mathbf{H}}^{m},\bar{\mathbf{y}}^{m}),$ (7)
where $f_{\Omega_{\rm DEC}}$ is the mapping function with variable parameters $\Omega_{\rm DEC}$.
@@ II-C Channel Prediction
In mobile scenarios involving high-velocity users, it is possible that the channel coherence time is shorter than the channel estimation period. In such cases, accurate channel prediction is vital to address the channel aging issue [20] in mobile communications with fast time-varying channels. In this work, we aim to accurately predict future channel state information (CSI) of $T_{2}$ time slots based on historical CSI of $T_{1}$ time slots. The CSI of $M$ subcarriers at time $t$ is represented in matrix form as
@@
The normalized MSE (NMSE) between predicted and actual future CSI is employed to evaluate the prediction accuracy. Then the entire problem can be described as follows:
@@
where $\hat{\mathbf{H}}^{t}$ is the predicted CSI, and $f_{\Omega_{\rm CP}}$ is the constructed mapping function and ${\Omega_{\rm CP}}$ is its variable parameters. Note here we ignore the subscript since the channel prediction is performed independently to each user.
III Proposed Physical Layer Multi-task LLM
In this section, we propose an LLM-empowered physical layer network to unify multiple tasks (channel prediction, multi-user precoding, as well as signal detection) with a single network. The proposed framework is illustrated in Fig. 1, including multi-task instruction module, input encoder, LLM backbone, and output decoder. We describe each component, together with the training schedule as follows.
Figure 1: The model framework of our method.
@@ III-A Multi-task Instruction Template
When training a single unified model for multiple different tasks, a basic and critical problem is how to distinguish each task. Fortunately, thanks to the remarkable text understanding capability, LLM allows users to input task requirements with natural language. Therefore, we design multi-task instruction template with task specific tokens as prompt to make each task easily distinguishable. Next, we introduce our multi-task instruction template in details.
@@
The first part is the task identifier token, the second part is the task and data description, and the third part is the instruction input. Task Identifier provides a distinct identifier for each task to reduce the ambiguity across various tasks. Task description attempts to input basic domain knowledge and illustrate the dataset for LLM to better understand the task. Take multi-user precoding as an example, the designed task description is presented by “Multi-user precoding aims to maximize the sum rate of multiple users through the design of base station precoder, to serve multiple users using a single base station. For the collected dataset, we consider a base station with 4 RF chains and 128 antennas to serve 4 single antenna user simutaneously.”. The third part Instruction gives direct and clear objective of the task. Again, for instance the instruction for multi-user precoding is “$<$Instruction$>$ Design the precoding matrix given channels of the users, to maximize the sum rate of the multiple users.”.
The designed prompt is then fed to the pre-trained LLM embedder to adapt to the input format of LLM. The prompt embedding is denoted as $\mathbf{X}_{\rm promot}^{\rm emb}$.

### v2

@@ Abstract
The advance of Artificial Intelligence (AI) is continuously reshaping the future 6G wireless communications. Particularly, the development of Large Language Models (LLMs) offers a promising approach to effectively improve the performance and generalization of AI in different physical-layer (PHY) tasks. However, most existing works finetune dedicated LLM networks for a single wireless communication task separately. Thus performing diverse PHY tasks requires extremely high training resources, memory usage, and deployment costs. To solve the problem, we propose a LLM-enabled multi-task PHY network to unify multiple tasks with a single LLM, by exploiting the excellent semantic understanding and generation capabilities of LLMs. Specifically, we first propose a multi-task LLM framework, which finetunes LLM to perform multi-user precoding, signal detection and channel prediction simultaneously. Besides, multi-task instruction module, input encoders, as well as output decoders, are elaborately designed to distinguish different tasks. The proposed design allows different wireless data types to be well aligned with the LLM input format. Moreover, low-rank adaptation (LoRA) is utilized for LLM fine-tuning. To reduce the memory requirement during LLM fine-tuning, a LoRA fine-tuning-aware quantization method is introduced. Extensive numerical simulations are also displayed to verify the effectiveness of the proposed method.
@@ I Introduction
The deep integration of Artificial Intelligence (AI) with wireless communications is one of the key features of Sixth-Generation (6G) communications [1]. With the development of 6G communications, the application of AI becomes essential to manage the explosively-growing data service. Attributed to the strong feature extraction capability, AI, especially deep learning (DL), has demonstrated great potential in a wide range of physical-layer (PHY) communication tasks [2], including CSI feedback [3], channel estimation [4], channel prediction [5], signal detection [6], etc. The AI Radio Access Network (AI-RAN) is anticipated to offer reduced latency, improved bandwidth, spectrum efficiency, and coverage. Moreover, significant efforts have been made by 3rd Generation Partnership Project (3GPP) to standardize DL in wireless networks recently [7].
@@
Despite the significant progress, existing DL empowered methods still face some fundamental issues that limit their applications in practical communication networks. First, with the dramatically increasing channel dimension and rapidly changing wireless environment in 6G applications, it is difficult for existing DL methods to comprehensively recognize patterns from complicated data distribution, due to their small model size and simple network structure. This insufficiency restricts their ability to provide reliable solutions in dynamic real-world scenarios. Secondly, existing DL methods exhibit a poor generalization capability to different wireless environments. For instance, a channel prediction model trained in an indoor scenario usually requires retraining when the wireless environment is changed to urban. Fortunately, recent advancements in large language models (LLMs) [8, 9, 10] have provided a radical solution to the challenges of existing DL methods. Bearing a huge amount of parameters, LLMs possess the ability to capture universal knowledge, which has demonstrated impressive language understanding and generation capabilities to various tasks in different domains.
Very recently, several initial studies [11, 12, 13, 14] have been conducted since 2024 to leverage LLMs for boosting the performance of PHY communication systems. To be specific, in [11] LLM is adopted to perform power allocation using a few-shot learning approach. In this method, the channel gains and the corresponding transmit power strategies are provided as the input of LLM, also called “prompt” in the society of AI. The paper [11] demonstrates that LLM can automatically understand the principle of water-filling based optimal power allocation without any retraining. The authors in [12] effectively utilize LLMs to enhance AI-based CSI feedback performance in various scenarios. They incorporate the channel distribution as a prompt within the decoder to further enhance channel reconstruction quality. Besides, in order to obtain more accurate channel prediction and improve the generalization capability, the authors in [13] propose a LLM-driven channel prediction approach.  [14] also unleashes the strength of LLMs for time series forecasting to improve the robustness of beam prediction. Moreover, a few review papers [15, 16, 17, 18, 19] have explored the potential transformative impact of LLMs and provide envisions for LLM-aided wireless communications.
However, most existing works finetune dedicated LLM networks for a single wireless communication task such as the previously mentioned channel prediction and CSI feedback, etc. In reality, wireless communication systems involve a multitude of tasks, each with distinct requirements that necessitate the adoption of LLMs. Given the huge model size of LLM, designing and training dedicated models for each task separately will lead to extremely high computational complexity, memory usage, and deployment costs.
To solve the problem, we propose a LLM-enabled multi-task PHY network to unify multiple tasks with a single LLM, by leveraging the excellent language understanding and generation capabilities of LLMs. The main contributions of this paper are summarized as follows [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html..
  • We propose a multi-task LLM framework for PHY communications. The framework enables us to input different task requirements to LLMs with natural language, and explore the global feature extraction ability of LLMs to execute multiple tasks within one network. Particularly, this work focuses on three of the most typical tasks in PHY communications, multi-user precoding, signal detection, and channel prediction. The proposed framework can also accommodate any other PHY tasks.
  • For the design of the proposed framework, dedicated modules are proposed to adapt the LLMs for multiple tasks in the wireless domain. It mainly consists of two components: multi-task instructions, and task-specific encoders and decoders. Specifically, first, in order to distinguish and cope with different tasks and different data formats of the tasks, we elaborately design multi-task instructions as prompts of LLM; secondly, to bridge the gap between the feature of wireless data and that of the LLM, task-specific encoders and decoders are proposed to adapt the text-based pre-trained LLM.
  • For the fine-tuning of the proposed framework, we introduce low-rank adaptation (LoRA) [20], which freezes the pre-trained model weights and injects trainable low-rank matrices. Furthermore, to mitigate the computational and memory demands of the proposed model, we employ a LoRA fine-tuning-aware quantization method [21] that simultaneously quantizes an LLM and finds a proper low-rank initialization for LoRA fine-tuning.
  • Extensive simulation experiments have been conducted to verify the effectiveness of the proposed method, which achieves comparable performance with the dedicated designed network for each task. In addition, the LoRA fine-tuning-aware quantization method achieves 75% storage reduction during fine-tuning with almost no performance degradation. Moreover the proposed multi-task instruction effectively accelerates the convergence and improve the performance.
The rest of the paper is organized as follows. Section II introduces the system model. Then the three selected tasks, namely multi-user precoding, signal detection, and channel prediction, are formulated respectively. In Section III, the design of the proposed multi-task PHY LLM is illustrated in detail. Besides, Section IV elaborates the fine-tuning strategy of the proposed multi-task PHY LLM. Simulation results are provided in Section V. Finally, Section VI concludes this paper.
@@ II Systems Model
We consider a multi-user (MU) multiple-input-single-output (MISO)- orthogonal-frequency-division-multiplexing (OFDM) system working in a time-division-duplex (TDD) mode. A base station (BS) simultaneously serves $K$ single-antenna mobile users using $K$ RF chains. The BS is equipped with a uniform planar array (UPA) consisting of $N_{\rm T}=N_{h}\times N_{v}$ antennas, where $N_{h}$ and $N_{v}$ denote the quantity of antennas along the horizontal and vertical dimensions, respectively. To design a multi-task LLM for the BS, we select three typical tasks in PHY communications for instance, downlink multi-user precoding, uplink signal detection, and channel prediction for mobile devices, respectively. The system models and problem descriptions of these three tasks are presented one by one below.
@@ II-A Multi-user Precoding
For downlink transmission scenario, the channel between user $k$ and the BS at the $m$-th subcarrier is denoted as ${{\bf h}_{k}^{m}}\in\mathbb{C}^{N_{t}\times 1},m=1,2,\cdots,M$. The received signal of user $k$ at the $m$-th subcarrier is given by
$y_{k}^{m}={\mathbf{h}_{k}^{m}}^{H}\sum_{k^{\prime}=1}^{K}\mathbf{w}_{k^{\prime}}^{m}x_{k^{\prime}}^{m}+n_{k}^{m},$ (1)
where $\mathbf{w}_{k}^{m}$ represents the beamforming vector for user $k$, $x_{k}^{m}$ with $\mathbb{E}(|x_{k}^{m}|^{2})=1$, is the transmitted symbol from the BS to user $k$, and $n_{k}^{m}\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive Gaussian white noise (AWGN) with zero mean and variance $\sigma^{2}$.
Multi-user precoding aims to maximize the system sum rate via the optimization of the transmit precoders. The total power of all beamforming vectors is limited due to the BS power budget. For simplicity, we design the beamformers based on the channel of the central carrier-frequency ${\bf h}_{k}$ and the problem is mathematically formulated as
$\max_{\mathbf{W}}\sum_{k=1}^{K}\log_{2}(1+\gamma_{k}),\quad\text{s.t. }\sum_{k=1}^{K}\|\mathbf{w}_{k}\|^{2}\leq P_{\max},$ (2)
where $\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\cdots,\mathbf{w}_{K}]$ is a set of beamforming vectors and $P_{\rm max}$ is the power budget. Besides, $\gamma_{k}$ represents the received signal-to-interference-plus-noise ratio (SINR) at user $k$. It is written as
As pointed out in [22], the optimal downlink beamforming vectors for (2) follow the structure as
where $p_{k}$ is the power allocated to the $k$-th use, and ${\lambda_{k}}$ is a positive parameter and $\sum_{k=1}^{K}{\lambda_{k}}=\sum_{k=1}^{K}p_{k}=P_{\rm max}$. The solution structure in (4) provides the required expert knowledge for the LLM-empowered beamforming design in (2). Given this knowledge, the LLM is only required to learn $2K$ key parameters $\mathbf{\lambda}=[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$ and $\mathbf{p}=[p_{1},p_{2},\cdots,p_{K}]$, instead of the whole $K\times N$ beamforming matrix $\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\cdots,\mathbf{w}_{K}]$. Thus, the problem of DL-based multi-user precoding can be reformulated as
@@
[…]
where $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{K}]$, and $f_{\Omega_{\rm PRE}}$ is the mapping function with variable parameters $\Omega_{\rm PRE}$.
@@ II-B Signal Detection
During uplink transmission, the BS antenna array simultaneously receive the transmitted symbols from the $K$ users. For clarity, we introduce the superscript “$\bar{\cdot}$” to all parameters related to the uplink transmission, e.g., $\bar{a}$. Specifically, we denote the transmitted symbol vector from all users at the $m$-th subcarrier as $\bar{\mathbf{x}}^{m}=[\bar{x}^{m}_{1},\bar{x}^{m}_{2},\cdots,\bar{x}^{m}_{K}]\in\mathbb{C}^{K\times 1}$. Each element is drawn from the $P$-QAM constellation, and then transmitted over channel. The received signal $\bar{\mathbf{y}}^{m}\in\mathbb{C}^{N_{\rm T}\times 1}$ is given by
@@
where $\bar{\mathbf{H}}^{m}=[\bar{\mathbf{h}}_{1}^{m},\bar{\mathbf{h}}_{2}^{m}.\cdots,\bar{\mathbf{h}}_{K}^{m}]$ is the uplink channel of the $K$ users and $\bar{\mathbf{n}}^{m}\sim\mathcal{CN}(0,\bar{\sigma}^{2}\mathbf{I}_{N_{\rm T}})$ is the AWGN at the $m$-th subcarrier.
BS requires to recover the signals $\bar{\mathbf{x}}^{m}$ from the received signal $\bar{\mathbf{y}}^{m}$ given $\bar{\mathbf{H}}^{m}$. We adopt the minimum mean squared error (MMSE) estimator to formulate the associated signal detection problem as
@@
s.t. $\hat{\bar{\mathbf{x}}}^{m}=f_{\Omega_{\rm DET}}(\bar{\mathbf{H}}^{m},\bar{\mathbf{y}}^{m}),$ (9)
where $f_{\Omega_{\rm DET}}$ is the mapping function with variable parameters $\Omega_{\rm DET}$.
@@ II-C Channel Prediction
In mobile scenarios involving high-velocity users, it is possible that the channel coherence time is shorter than the channel estimation period. Under these circumstances, precise channel prediction becomes crucial to mitigate the channel aging phenomenon [23] in high-mobility communication environments with rapidly changing channels. In this work, we aim to accurately predict channel state information (CSI) of next $T_{2}$ time slots given CSI of previous $T_{1}$ time slots. The CSI of $M$ subcarriers at time $t$ is represented in matrix form as
@@
To evaluate the channel prediction accuracy, the normalized MSE (NMSE) between predicted CSI by the network and ground-truth CSI is selected as the metric. Utilizing the metric, the channel prediction problem can be described as follows:
@@
where $\hat{\mathbf{H}}^{t}$ is the predicted CSI, and $f_{\Omega_{\rm CP}}$ is the mapping function of the network with trainable parameters ${\Omega_{\rm CP}}$. Note here we ignore the subscript since the channel prediction is performed independently to each user.
III The Framework of the Proposed Physical Layer Multi-task LLM
In this paper, we propose an LLM-enabled PHY network to unify multiple tasks (channel prediction, multi-user precoding, as well as signal detection) with a single network. The proposed framework is illustrated in Fig. 1, including a multi-task instruction module, an input encoder, an LLM backbone, and an output decoder. To distinguish and cope with different tasks, we design multi-task instructions as prompts of LLM, which are processed by the pre-trained LLM embedder as part of the LLM input. Besides, the wireless data are encoded by task-specific encoders to make the feature of wireless data understandable to a text-based pre-trained LLM, e.g., GPT-2. The prompt and encoded wireless data are concatenated together to serve as the inputs of LLM. These inputs are later fed into the LLM backbone. Note that the same LLM backbone is used for all considered tasks. Finally, obtaining the outputs of the LLM backbone, task-specific decoders are utilized to generate desired outputs for different tasks.
We describe each component of the framework as follows. It is worth noting that, in this paper, we select three typical tasks suitable for deep learning methods in PHY communications, namely multi-user precoding, signal detection, and channel prediction; while the proposed method can be smoothly extended to other PHY tasks. Besides, we mention that although the encoders and decoders are designed specifically for each task, the parameter sizes of these modules are much smaller than that of the LLM backbone. The LLM backbone, which takes up the majority of the total parameter size, is shared among different tasks.
Figure 1: The model framework of our method. In this figure, “CP” denotes channel prediction, “Det” denotes signal detection, and “Pre” denotes multi-user precoding.
@@ III-A Multi-task Instruction Template
When training a single unified model for multiple different tasks, a basic and critical problem is how to distinguish each task. Thanks to the remarkable text understanding capability, LLM allows users to input task requirements with natural language. Therefore, we can design a multi-task instruction template with task-specific tokens as prompt to make each task easily distinguishable. Next, we introduce our multi-task instruction template in detail.
@@
The first part is the task identifier token, the second part is the task and data description, and the third part is the instruction input. Task Identifier provides a distinct identifier for each task to reduce the ambiguity across various tasks. Task description attempts to input basic domain knowledge and illustrate the dataset for LLM to better understand the task. Take multi-user precoding as an example, the designed task description is presented by “Multi-user precoding aims to maximize the sum rate of multiple users through the design of precoder. For the collected dataset, we consider a base station with 128 antennas to serve 8 single antenna users simultaneously”. The third part Instruction gives direct and clear objective of the task. Again, for instance the instruction for multi-user precoding is “$<$Instruction$>$ Design the precoding matrix given channels of the users, to maximize the sum rate of the multiple users.”.
The designed prompt is then fed to the pre-trained LLM embedder as part of the input of LLM. The prompt embedding is denoted as $\mathbf{X}_{\rm promot}^{\rm emb}$.